OpenStudy (anonymous):

The height of a cylinder with a fixed radius of 10 cm is increasing at the rate of 0.5 cm/min. Find the rate of change of the volume of the cylinder (with respect to time) when the height is 30cm.

OpenStudy (anonymous):

this link would help u buddy https://www.jackmathsolutions.com/images/CHAPTER_8.pdf

OpenStudy (dls):

Do you know the formula for volume of cylinder?

OpenStudy (dls):

\[\LARGE V= \pi r^2 H\] Differentiate wrt to time \[\LARGE \frac{dV}{dt} = \pi r^2 \times \frac{dh}{dt}\] Note that r is given constant here. Use the info from your question that: r=10 dr/dt= 0.5 h=30 Substitute everything and solve for LHS.

OpenStudy (anonymous):

Thanks, I already got, actually finding that link on my own @gowtham89 lol